# The Tarantula Massive Binary Monitoring

### v1

@@ Appendix D Spectral disentangling
In case of our IFU observations of VFTS 812, the imperfect nebular line subtraction led to residuals in the sky-subtracted spectra that can considerably affect the disentangling and introduce artifacts in the disentangled spectra. To mitigate this effect, we change to a vector quantity defined for all . We retain the continuum noise value for most wavelength bins, but arbitrarily increase to 999 for , which is affected by the nebular line as discussed in Appendix B. With this modification, we can perform a meaningful spectral disentangling analysis despite residuals from sky subtraction.
D.1 Input parameters
@@
D.2 Component flux ratios
For the injection study discussed in Section 4.2, the mock data need to be synthesized using appropriate flux ratios between the primary and the injected secondary. Similar to Shenar et al. (2022b), we convolve our flux calibrated fastwind models with the Bessel B filter and take the ratio subsequently. Table 5 lists the computed flux ratios between different secondary models and the primary model.
@@ D.2 Component flux ratios
Appendix E Photometric constraints
In Section 3.3, the lack of observed photometric variability in VFTS 812 was used to obtain an upper limit on the inclination, on the basis that a more edge-on orbit would produce an observable primary eclipse. Figure E.1 shows the OGLE I-band light curve that TS22 remarked on for not having any significant frequencies. Using a 3- threshold, we find that the flux-dip from an eclipse has to be 2.6% for detection. From spectral disentangling, we know that the flux contribution of the secondary is 1.6% (see Sec. 4.2). A full eclipse of the primary by a companion with a radius of will cover 5% of the primary surface, while contributing 1.6% to the total flux, thus resulting in a flux-dip of 3.4%. While partial eclipses can be considered for finer estimation of eclipse detectability, we adopt the full eclipse scenario to remain conservative. The lack of a detected eclipse therefore enables an additional inclination constraint on the MS companion scenario.

### v2

@@ Appendix F Spectral disentangling
In case of our IFU observations of VFTS 812, the imperfect nebular line subtraction led to residuals in the sky-subtracted spectra that can considerably affect the disentangling and introduce artifacts in the disentangled spectra. To mitigate this effect, we change to a vector quantity defined for all . We retain the continuum noise value for most wavelength bins, but arbitrarily increase to 999 for , which is affected by the nebular line as discussed in Appendix C. With this modification, we can perform a meaningful spectral disentangling analysis despite residuals from sky subtraction.
F.1 Input parameters
@@
F.2 Component flux ratios
For the injection study discussed in Sect. 4.2, the mock data need to be synthesized using appropriate flux ratios between the primary and the injected secondary. Similar to Shenar et al. (2022b), we convolve our flux calibrated fastwind models with the Bessel B filter and take the ratio subsequently. Table 5 lists the computed flux ratios between different secondary models and the primary model.